Circular No. 7437

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cm IN ANONYMOUS GALAXY S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrogram of SN 2000cm (cf. IAUC 7436), obtained by P. Berlind on June 4.16 and 5.17 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ia supernova after maximum light. The age of the supernova based on its spectral features (Riess et al. 1997, A.J. 114, 722) is 14 ± 3 days after maximum light. Cross-correlation of the host-galaxy spectrum with emission and absorption-line templates yields a recession velocity of 2170 ± 30 km/s.

SUPERNOVA 2000cn IN UGC 11064 Jha et al. report that a spectrum of SN 2000cn (cf. IAUC 7436), taken

by M. Calkins on June 3.37 UT with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ia supernova well before maximum light. Adopting the NASA/IPAC Extragalactic Database rescession velocity for the host galaxy (7043 km/s), the supernova expansion velocities measured with the Ca H and K and the Si II (rest 635.5 nm) features are 17 400 km/s and 12 900 km/s, respectively. For typical type-Ia supernovae, these high velocities imply an age of ∼ 7 to 10 days before maximum light (Jha et al. 1999, Ap.J. Supp. 125, 73).

M. Turatto, A. Pastorello, E. Cappellaro, and F. Cedrati, Osservatorio Astronomico di Padova, write: “We obtained spectra (range 350–750 nm, resolution 2 nm) on June 5.0 UT with the Asiago 1.82-m telescope (+ AFOSC), showing SN 2000cn to be of type Ia, but very close to maximum. Si II 635-nm is measured at 624.0 nm. This, combined with the known galaxy recession velocity (7050 km/s as reported in NED), gives a photospheric expansion velocity of 12 500 km/s. Other strong absorptions are measured at 384, 440, 496, and 587 nm.”

SUPERNOVA 2000co IN MCG +7-33-17 Jha et al. also report that a spectrum of SN 2000co (cf. IAUC 7436),

taken by Berlind on June 4.28 UT as above, shows it to be a type-II supernova at maximum light. The spectrum exhibits a flat continuum with broad P-Cyg Balmer lines and He I. Cross-correlation of the host-galaxy spectrum with an emission-line template yields a recession velocity of 11 590 ± 30 km/s. The expansion velocity of the supernova, measured via the Hα profile, is 7300 km/s.

2000 June 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green